OpenStudy (anonymous):

3. Prove that any open ball in a metric space (X, d) is open and any closed ball is closed.

OpenStudy (zzr0ck3r):

Let \(X\) be a metric space and let \(G=B(x_0;r)\) be an open ball in \(X\). Fix any \(a\in G\). Then \(d(x_0,a)<r\), thus \(\delta = r - d(x_0,a)>0.\) Let \(x\in B(a;\delta)\), then \(d(x,a)< \delta = r-d(x_0,a)\). It follows that \(d(x,x_0)\le d(x,a)+d(a,x_0)<r-d(x_0,a)+d(x_0,a)=r\) as desired, and thus \(B(x_0,r)\) is an open set.

OpenStudy (zzr0ck3r):

Let me know if this makes sense then we can move on to the closed ball problem.

OpenStudy (anonymous):

yes.. i think that makes sense..let me try it first..